# Mitigating the Modality Gap: Few-Shot Out-of-Distribution Detection with Multi-modal Prototypes and Image Bias Estimation

### v1

@@ 7.2 Ablations
Evaluating the effectiveness of the proposed modules (BPG and ITC) and $\boldsymbol{S_{\textit{GMP}}}$ (complete results). To evaluate the effectiveness of the proposed modules and $S_{\textit{GMP}}$, we present detailed results of a combinatorial ablation study in Tab. 9. Performance is evaluated across four datasets (i.e., iNaturalist, SUN, Places, and Texture).
Compared with the base model (MCM [45]), introducing ITC considerably reduces FPR95 to 28.29 while boosting AUROC to 93.40. Similarly, BPG improves performance, though to a slightly lesser extent, achieving an average FPR95 of 29.89 and AUROC of 93.07. When ITC and BPG are combined, the performance improves further, reaching an average FPR95 of 25.03 and AUROC of 94.80. These results show that both ITC and BPG individually contribute substantial improvements over the base model, while their combination yields even greater performance.
For the proposed score, $S_{\textit{GMP}}$, when combining it with either ITC or BPG, the performance shows incremental gains compared to using ITC or BPG alone. The best results are obtained when all three components (ITC, BPG, and $S_{\textit{GMP}}$) are combined, yielding an average FPR95 of 20.28 and AUROC of 95.54. This configuration achieves the lowest FPR95 and highest AUROC across all configurations and datasets.
Combining suPreMe with post-hoc methods. To understand how suPreMe combines with existing representative post-hoc methods, i.e., ReAct [55] and $S_{\textit{EOE}}$ [4], we present a systematic study in Tab. 9. We directly apply ReAct and $S_{\textit{EOE}}$ on our best model as post-hoc methods, which means we do not use them during training. For ReAct, we set the threshold at 0.95, while for $S_{\textit{EOE}}$, we follow the standard parameters. We see that using ReAct alone makes the performance worse with average FPR95 of 21.21 and AUROC of 94.35. On the other hand, $S_{\textit{EOE}}$ boosts average FPR95 from 20.28 to 19.25. When combining ReAct and $S_{\textit{EOE}}$, the FPR95 is boosted to 19.14 on average. However, using ReAct and/or $S_{\textit{EOE}}$ decreases the average AUROC compared to vanilla suPreMe, showing that these post-hoc methods might only help reduce false positives. The experiments show that suPreMe is compatible with these post-hoc methods without performance drops.
@@
(e) $S_{\textit{GMP}},\textit{KS}=0.8724$.
(f) $S_{\textit{GMP}},\textit{KS}=0.7798$.
(g) $S_{\textit{GMP}},\textit{KS}=0.7219$.
(h) $S_{\textit{GMP}},\textit{KS}=0.6874$.
@@
How much data do we need for few-shot tuning? We evaluate the effect of the number of shots for few-shot tuning on the model’s performance in Fig. 6. As the number of shots increases from $1$ to $16$, we observe a consistent improvement in both metrics. Specifically, FPR95 decreases from 28.73 at $1$-shot to 20.28 at $16$-shot, indicating that more shots lead to better separability between ID and OOD data. Similarly, AUROC improves from 92.13 at $1$-shot to a peak of 95.54 at $16$-shot, demonstrating enhanced overall detection performance. These results highlight the robustness of suPreMe in leveraging additional labeled examples, with diminishing returns on FPR95 as data size increases. While higher shot numbers provide the best performance, the model performs reasonably well even with a limited number of shots, showcasing its flexibility in low-data scenarios.
@@
Hyperparameter ($\alpha$ and $\beta$) sensitivity. We analyze the effect of the trade-off hyperparameters $\alpha$ and $\beta$ (Sec. 4.2) on the model’s performance, as shown in Fig. 7. For $\alpha$, we observe a trade-off between FPR95 and AUROC as its value increases. Specifically, as $\alpha$ grows from 0.001 to 0.01, FPR95 decreases initially from 20.82 to 20.28 at $\alpha=0.005$, indicating improved separability, but then rises to 25.56 at $\alpha=0.01$. Meanwhile, AUROC peaks at 95.54 for $\alpha=0.005$ before declining, indicating that excessively large values of $\alpha$ may overemphasize certain features, leading to diminished overall performance. Similarly, $\beta$ demonstrates a comparable trend. The FPR95 achieves its lowest value of 20.28 at $\beta=0.1$, while AUROC peaks at 95.54. The results of both hyperparameters suggest that suPreMe is not sensitive to the parameters and that moderate values ($\alpha=0.005,\beta=0.1$) balance the trade-off between false positive rate and overall discriminative ability.
@@
Full statistical evaluation of $\boldsymbol{S_{\textit{GMP}}}$. We visualize the empirical cumulative distribution (empirical CDF) and the Kolmogorov–Smirnov (KS) test statistics for $S_{\textit{MCM}}$ and our proposed $S_{\textit{GMP}}$ in Fig. 8. The evaluation uses our best model trained on the ImageNet-1k (ID) dataset, with iNaturalist, SUN, Places, and Texture serving as OOD datasets. The empirical CDF clearly shows that $S_{\textit{GMP}}$ improves the gap between ID and OOD data more effectively. Furthermore, the KS test statistics improve considerably from $0.7270,0.7182,0.6650,$ and $0.6760$ ($S_{\textit{MCM}}$) to $0.8724,0.7798,0.7219,$ and $0.6874$ ($S_{\textit{GMP}}$) across the four OOD datasets. These results indicate that $S_{\textit{GMP}}$ achieves better separation between ID and OOD scores compared to $S_{\textit{MCM}}$.
@@ 7.3 Details of Benchmarking Datasets
  • iNaturalist contains images from the natural world with images from 5089 classes, belonging to 13 super-categories, such as Plantae (Plant), Insecta (Insect), Aves (Bird), Mammalia (Mammal). The subset containing 110 plant classes not present in ImageNet-1k is chosen as the OOD test set. The classes are as follows, Coprosma lucida, Cucurbita foetidissima, Mitella diphylla, Selaginella bigelovii, Toxicodendron vernix, Rumex obtusifolius, Ceratophyllum demersum, Streptopus amplexifolius, Portulaca oleracea, Cynodon dactylon, Agave lechuguilla, Pennantia corymbosa, Sapindus saponaria, Prunus serotina, Chondracanthus exasperatus, Sambucus racemosa, Polypodium vulgare, Rhus integrifolia, Woodwardia areolata, Epifagus virginiana, Rubus idaeus, Croton setiger, Mammillaria dioica, Opuntia littoralis, Cercis canadensis, Psidium guajava, Asclepias exaltata, Linaria purpurea, Ferocactus wislizeni, Briza minor, Arbutus menziesii, Corylus americana, Pleopeltis polypodioides, Myoporum laetum, Persea americana, Avena fatua, Blechnum discolor, Physocarpus capitatus, Ungnadia speciosa, Cercocarpus betuloides, Arisaema dracontium, Juniperus californica, Euphorbia prostrata, Leptopteris hymenophylloides, Arum italicum, Raphanus sativus, Myrsine australis, Lupinus stiversii, Pinus echinata, Geum macrophyllum, Ripogonum scandens, Echinocereus triglochidiatus, Cupressus macrocarpa, Ulmus crassifolia, Phormium tenax, Aptenia cordifolia, Osmunda claytoniana, Datura wrightii, Solanum rostratum, Viola adunca, Toxicodendron diversilobum, Viola sororia, Uropappus lindleyi, Veronica chamaedrys, Adenocaulon bicolor, Clintonia uniflora, Cirsium scariosum, Arum maculatum, Taraxacum officinale officinale, Orthilia secunda, Eryngium yuccifolium, Diodia virginiana, Cuscuta gronovii, Sisyrinchium montanum, Lotus corniculatus, Lamium purpureum, Ranunculus repens, Hirschfeldia incana, Phlox divaricata laphamii, Lilium martagon, Clarkia purpurea, Hibiscus moscheutos, Polanisia dodecandra, Fallugia paradoxa, Oenothera rosea, Proboscidea louisianica, Packera glabella, Impatiens parviflora, Glaucium flavum, Cirsium andersonii, Heliopsis helianthoides, Hesperis matronalis, Callirhoe pedata, Crocosmia × crocosmiiflora, Calochortus albus, Nuttallanthus canadensis, Argemone albiflora, Eriogonum fasciculatum, Pyrrhopappus pauciflorus, Zantedeschia aethiopica, Melilotus officinalis, Peritoma arborea, Sisyrinchium bellum, Lobelia siphilitica, Sorghastrum nutans, Typha domingensis, Rubus laciniatus, Dichelostemma congestum, Chimaphila maculata, Echinocactus texensis.
@@ 8 Proof of Theorem 1
Assumption 4 (OOD embeddings are not sampled from any ID distribution.).
OOD image embeddings $\mathbf{x}_{\textit{OOD}}$ are drawn from a distribution $\mathcal{D}_{\text{{OOD}}}$ different from any $\mathcal{D}_{c}^{*},\forall c\in[C],*\in[\textit{text},\textit{image}]$. Besides, the OOD image embeddings are equally similar to all ID text/image classes.
@@
This assumption is a general basic assumption on the OOD data, which provides us with an overall characteristic of the OOD data.
Under the above assumptions, we first would like to prove the following inequality holds,
$\Delta_{\text{image}}=\mathbb{E}\left[\max_{i\in[C]}\sigma_{i}^{\text{image}}(\mathbf{x}_{\textit{ID}})-\max_{i\in[C]}\sigma_{i}^{\text{image}}(\mathbf{x}_{\textit{OOD}})\right]$ (12)
$\geq$ $\Delta_{\text{text}}=\mathbb{E}\left[\max_{i\in[C]}\sigma_{i}^{\text{text}}(\mathbf{x}_{\textit{ID}})-\max_{i\in[C]}\sigma_{i}^{\text{text}}(\mathbf{x}_{\textit{OOD}})\right]\,,$
[…]
@@
$\mathbb{E}\left[\max_{c\in[C]}\sigma_{c}^{*}(\mathbf{x}_{\textit{ID}})\right]$ $=\frac{\exp\left(\mu_{\text{intra}}^{*}\right)}{\exp\left(\mu_{\text{intra}}^{*}\right)+(C-1)\exp\left(\mu_{\text{inter}}^{*}\right)},$ (13)
[…]
where $\mu_{\textit{intra}}^{*}=\max_{c\in[C]}\mathbb{E}[cos(\mathbf{x}_{\textit{ID}},P_{*,c})]$, $\mu_{iter}^{*}=\mathbb{E}[cos(\mathbf{x}_{\textit{ID}},I_{c_{j}})],\forall c_{j}\in[C]\setminus\{c_{\textit{ID}}\}$, and $C_{\textit{ID}}$ is the ground-truth class for the input $\mathbf{x}_{\textit{ID}}$.
With assumption 4, for OOD samples, we have,
$\mathbb{E}\left[\max_{c\in[C]}\sigma_{c}^{\text{image}}(\mathbf{x}_{\textit{OOD}})\right]=\mathbb{E}\left[\max_{c\in[C]}\sigma_{c}^{\text{text}}(\mathbf{x}_{\textit{OOD}})\right]=\frac{1}{C}\,.$ (14)
Then we have,
$\Delta_{\text{image}}=\mathbb{E}\left[\max_{c\in[C]}\sigma_{c}^{\text{image}}(\mathbf{x}_{i})\right]-\frac{1}{C}\,.$ (15)
Similarly, we have,
$\Delta_{\text{text}}=\mathbb{E}\left[\max_{c\in[C]}\sigma_{c}^{\text{text}}(\mathbf{x}_{i})\right]-\frac{1}{C}\,.$ (16)
Since $\mu_{\text{intra}}^{\text{image}}>\mu_{\text{intra}}^{\text{text}}$ (Assumption 2), we have,
$\Delta_{\text{image}}$ $>\Delta_{\text{text}}\,.$ (17)
By simple algebra, we have,
$\mathbb{E}\left[\frac{\max_{i\in[C]}\sigma_{i}^{\text{image}}(\mathbf{x}_{\textit{ID}})+\max_{i\in[C]}\sigma_{i}^{\text{text}}(\mathbf{x}_{\textit{ID}})}{2}\right.$ (18)
$\left.-\frac{\max_{i\in[C]}\sigma_{i}^{\text{image}}(\mathbf{x}_{\textit{OOD}})+\max_{i\in[C]}\sigma_{i}^{\text{text}}(\mathbf{x}_{\textit{OOD}})}{2}\right]$
$\geq$ $\mathbb{E}\left[\max_{i\in[C]}\sigma_{i}^{\text{text}}(\mathbf{x}_{\textit{ID}})-\max_{i\in[C]}\sigma_{i}^{\text{text}}(\mathbf{x}_{\textit{OOD}})\right]\,,$

### v2

@@ 9.2 Ablations
Evaluating the effectiveness of the proposed modules (BPG and ITC) and $\boldsymbol{S_{\textit{GMP}}}$ (full results). To evaluate the effectiveness of the proposed modules and $S_{\textit{GMP}}$, we present detailed results of a combinatorial ablation study in Tab. 11. Performance is evaluated across four datasets (i.e., iNaturalist, SUN, Places, and Texture).
Combining suPreMe with post-hoc methods. To understand how suPreMe combines with existing representative post-hoc methods, i.e., ReAct [53] and $S_{\textit{EOE}}$ [3], we present a systematic study in Tab. 12. We directly apply ReAct and $S_{\textit{EOE}}$ on our best model as post-hoc methods, which means we do not use them during training. For ReAct, we set the threshold at 0.95, while for $S_{\textit{EOE}}$, we follow the standard parameters. We see that using ReAct alone makes the performance worse with an average FPR95 of 20.82 and AUROC of 94.72. On the other hand, $S_{\textit{EOE}}$ boosts average FPR95 from 20.70 to 19.63. When combining ReAct and $S_{\textit{EOE}}$, the FPR95 is boosted to 19.55 on average. However, using ReAct and/or $S_{\textit{EOE}}$ decreases the average AUROC compared to vanilla suPreMe, showing that these post-hoc methods might only help reduce false positives. The experiments show that suPreMe is compatible with these post-hoc methods without performance drops.
Effectiveness of suPreMe with different backbones. The results in Tab. 13 show that suPreMe consistently outperforms other methods across different backbones, including ResNet-50 [16] and ResNet-101 [16]. Specifically, suPreMe achieves 30.55 for FPR95 and 92.90 for AUROC with ResNet-50, and 25.30 for FPR95 and 93.81 for AUROC with ResNet-101, demonstrating its robustness and effectiveness across various model architectures.
@@
(e) $S_{\textit{GMP}},\textit{KS}=0.8555$.
(f) $S_{\textit{GMP}},\textit{KS}=0.7732$.
(g) $S_{\textit{GMP}},\textit{KS}=0.7281$.
(h) $S_{\textit{GMP}},\textit{KS}=0.6848$.
@@
$\left\|\overrightarrow{\Delta}_{\text{gap}}\right\|$ Modality 1 Modality 2 ImageNet (ID) $\downarrow$ iNaturalist $\downarrow$ Places $\downarrow$ Texture $\downarrow$ SUN $\downarrow$
MCM $P_{\text{txt}}$} II 0.8679 0.9839 0.9101 0.9373 0.9252
suPreMe $P_{\text{txt}}$} II 0.6135 0.7159 0.8961 0.9273 0.9002
$P_{\text{txt}}$} $f_{\text{img-txt}}(I)$ 0.1938 0.2238 0.2376 0.2559 0.2454
$P_{\text{img}}$} II 0.1348 0.4531 0.3474 0.3333 0.3447
Table 15: Quantifying the modality gap with different methods. We use CLIP (ViT-B/16) as the base model.
Impact of the length of prompts. In the main experiments, we follow previous works [45, 73, 74] and set the length of prompts to $L=16$. Here, we would like to validate that $16$ is the best choice for suPreMe. We conducted experiments with suPreMe and results shown in Fig. 7 indicate that among the choices from $1$ to $16$, $16$ obtains the lowest FPR95 and highest AUROC. Increasing the length of prompts improves the OOD detection performance.
@@
How much data do we need for few-shot tuning? We evaluate the effect of the number of shots for few-shot tuning on the model’s performance in Fig. 8 with suPreMe . As the number of shots increases from $1$ to $16$, we observe a consistent improvement in both metrics. Specifically, FPR95 decreases from 28.73 at $1$-shot to 20.28 at $16$-shot, indicating that more shots lead to better separability between ID and OOD data. Similarly, AUROC improves from 92.13 at $1$-shot to a peak of 95.54 at $16$-shot, demonstrating enhanced overall detection performance. These results highlight the robustness of suPreMe in leveraging additional labeled examples, with diminishing returns on FPR95 as data size increases. While higher shot numbers provide the best performance, the model performs reasonably well even with a limited number of shots, showcasing its flexibility in low-data scenarios.
Impact of base image size for constructing ID image prototypes. In $S_{\textit{GMP}}$, we use a set of base images to construct image prototypes $\{P_{i,c}\}_{c=1}^{C}$. To understand how the size of base images affects performance, we conduct experiments varying its size by randomly selecting a subset of the training images. The results in Fig. 9 show that by increasing the number of images for each class, the performance of suPreMe is improved, as image prototypes are closer to the real distribution. FPR95 and AUROC reach the best when using all the images. However, as shown in Tab. 2, the performance is already competitive when using the same set of few-shot data for fine-tuning, indicating that the model is robust to the size of base images.
@@
Hyperparameter ($\alpha$ and $\beta$) sensitivity. We analyze the effect of the trade-off hyperparameters $\alpha$ and $\beta$ (Sec. 4.2) on the model’s performance, as shown in Fig. 10 with suPreMe . For $\alpha$, we observe a trade-off between FPR95 and AUROC as its value increases. Specifically, as $\alpha$ grows from 0.001 to 0.01, FPR95 decreases initially from 20.82 to 20.28 at $\alpha=0.005$, indicating improved separability, but then rises to 25.56 at $\alpha=0.01$. Meanwhile, AUROC peaks at 95.54 for $\alpha=0.005$ before declining, indicating that excessively large values of $\alpha$ may overemphasize certain features, leading to diminished overall performance. Similarly, $\beta$ demonstrates a comparable trend. The FPR95 achieves its lowest value of 20.28 at $\beta=0.1$, while AUROC peaks at 95.54. The results of both hyperparameters suggest that suPreMe is not sensitive to the parameters and that moderate values ($\alpha=0.005,\beta=0.1$) balance the trade-off between false positive rate and overall discriminative ability.
@@
Full statistical evaluation of $\boldsymbol{S_{\textit{GMP}}}$. We visualize the empirical cumulative distribution (empirical CDF) and the Kolmogorov–Smirnov (KS) test statistics for $S_{\textit{MCM}}$ and our proposed $S_{\textit{GMP}}$ in Fig. 11. The evaluation uses our best model trained on the ImageNet-1k (ID) dataset, with iNaturalist, SUN, Places, and Texture serving as OOD datasets. The empirical CDF clearly shows that $S_{\textit{GMP}}$ improves the gap between ID and OOD data more effectively. Furthermore, the KS test statistics improve ceonsiderably from $0.7270,0.7182,0.6650,$ and $0.6760$ ($S_{\textit{MCM}}$) to $0.8724,0.7798,0.7219,$ and $0.6874$ ($S_{\textit{GMP}}$) across the four OOD datasets. These results indicate that $S_{\textit{GMP}}$ achieves better separation between ID and OOD scores compared to $S_{\textit{MCM}}$.
Quantifying modality gap. To quantify the modality gap, following [39], we formulate the modality gap as the $\ell_{2}$ distance between the center of image embeddings and (class) text embeddings, i.e., $\left\|\overrightarrow{\Delta}_{\text{gap}}\right\|=\left\|\frac{1}{C}\sum_{c=1}^{C}P_{\text{txt},c}-\frac{1}{N_{\text{image}}}\sum_{i=1}^{N_{\text{image}}}I_{i}\right\|_{2}$, where $P_{\text{txt},c}$ is the embedding of $c$-th class, $N_{\text{image}}$ is the number of images, and $I_{i}$ is the image embedding of the $i$-th image. We use the test set of each dataset for calculating the modality gap. We also calculate the modality gap between the center of mapped image embeddings $f_{\textit{txt-img}}(I)$ and image ID prototypes $P_{\text{img}}$} as shown in Tab. 15. The results demonstrate that suPreMe significantly reduces the modality gap across all datasets. Specifically, compared to the baseline MCM, suPreMe reduces the original modality gap by approximately 25-30%, which shows that the design of suPreMe could effectively reduce the gap. Moreover, when comparing text prototypes with our mapped image embeddings $f_{t2i}(I)$, we achieve a dramatic reduction of over 75% in the modality gap (row 3), indicating that our mapping function effectively bridges the gap between modalities. Additionally, the comparison between image ID prototypes and original images (row 4) shows substantial improvement.
@@ 9.3 Details of Benchmarking Datasets
  • iNaturalist contains images from the natural world with images from 5089 classes, belonging to 13 super-categories, such as Plantae (Plant), Insecta (Insect), Aves (Bird), Mammalia (Mammal). The subset containing 110 plant classes not present in ImageNet-1k is chosen as the OOD test set. The classes are as follows: Coprosma lucida, Cucurbita foetidissima, Mitella diphylla, Selaginella bigelovii, Toxicodendron vernix, Rumex obtusifolius, Ceratophyllum demersum, Streptopus amplexifolius, Portulaca oleracea, Cynodon dactylon, Agave lechuguilla, Pennantia corymbosa, Sapindus saponaria, Prunus serotina, Chondracanthus exasperatus, Sambucus racemosa, Polypodium vulgare, Rhus integrifolia, Woodwardia areolata, Epifagus virginiana, Rubus idaeus, Croton setiger, Mammillaria dioica, Opuntia littoralis, Cercis canadensis, Psidium guajava, Asclepias exaltata, Linaria purpurea, Ferocactus wislizeni, Briza minor, Arbutus menziesii, Corylus americana, Pleopeltis polypodioides, Myoporum laetum, Persea americana, Avena fatua, Blechnum discolor, Physocarpus capitatus, Ungnadia speciosa, Cercocarpus betuloides, Arisaema dracontium, Juniperus californica, Euphorbia prostrata, Leptopteris hymenophylloides, Arum italicum, Raphanus sativus, Myrsine australis, Lupinus stiversii, Pinus echinata, Geum macrophyllum, Ripogonum scandens, Echinocereus triglochidiatus, Cupressus macrocarpa, Ulmus crassifolia, Phormium tenax, Aptenia cordifolia, Osmunda claytoniana, Datura wrightii, Solanum rostratum, Viola adunca, Toxicodendron diversilobum, Viola sororia, Uropappus lindleyi, Veronica chamaedrys, Adenocaulon bicolor, Clintonia uniflora, Cirsium scariosum, Arum maculatum, Taraxacum officinale officinale, Orthilia secunda, Eryngium yuccifolium, Diodia virginiana, Cuscuta gronovii, Sisyrinchium montanum, Lotus corniculatus, Lamium purpureum, Ranunculus repens, Hirschfeldia incana, Phlox divaricata laphamii, Lilium martagon, Clarkia purpurea, Hibiscus moscheutos, Polanisia dodecandra, Fallugia paradoxa, Oenothera rosea, Proboscidea louisianica, Packera glabella, Impatiens parviflora, Glaucium flavum, Cirsium andersonii, Heliopsis helianthoides, Hesperis matronalis, Callirhoe pedata, Crocosmia × crocosmiiflora, Calochortus albus, Nuttallanthus canadensis, Argemone albiflora, Eriogonum fasciculatum, Pyrrhopappus pauciflorus, Zantedeschia aethiopica, Melilotus officinalis, Peritoma arborea, Sisyrinchium bellum, Lobelia siphilitica, Sorghastrum nutans, Typha domingensis, Rubus laciniatus, Dichelostemma congestum, Chimaphila maculata, Echinocactus texensis.
@@ 10 Proof of Theorem 1
Assumption 4 (OOD embeddings are not sampled from any ID distribution [12]).
OOD image embeddings $\mathbf{x}_{\textit{OOD}}$ are drawn from a distribution $\mathcal{D}_{\textit{OOD}}$ different from any $\mathcal{D}_{c}^{*},\forall c\in\{1..C\}\text{ and }\forall*\in\{\textit{text},\textit{image}\}$. In addition, the OOD image embeddings are equally similar to all ID text/image classes.
@@
This assumption is a general basic assumption used by previous studies [12] on the OOD data, which provides us with an overall characteristic of the OOD data.
Under the above assumptions, we first would like to prove the following inequality holds:
$\Delta_{\textit{image}}$ $=\mathbb{E}\left[\max_{i\in\{1..C\}}\sigma_{i}^{\textit{image}}(\mathbf{x}_{\textit{ID}})-\max_{i\in\{1..C\}}\sigma_{i}^{\textit{image}}(\mathbf{x}_{\textit{OOD}})\right]$ (12)
$\geq\Delta_{\textit{text}}$ $=\mathbb{E}\left[\max_{i\in\{1..C\}}\sigma_{i}^{\textit{text}}(\mathbf{x}_{\textit{ID}})-\max_{i\in\{1..C\}}\sigma_{i}^{\textit{text}}(\mathbf{x}_{\textit{OOD}})\right]$
$\text{s.t. }\sigma_{i}^{*}=\frac{\exp(s_{*,i})}{\sum_{k=1}^{N}\exp(s_{k})}\text{ and }s_{*,i}=cos(\mathbf{x},P_{*,i})$
$\ \ \forall*\in\{\textit{image},\textit{text}\}\,$,
@@
$\mathbb{E}\left[\max_{c\in\{1..C\}}\sigma_{c}^{*}(\mathbf{x}_{\textit{ID}})\right]$ $=\frac{\exp\left(\mu_{\textit{intra}}^{*}\right)}{\exp\left(\mu_{\textit{intra}}^{*}\right)+(C-1)\exp\left(\mu_{\textit{inter}}^{*}\right)}$ (13)
$\forall*\in\{\textit{text},\textit{image}\}\,$,
where $\mu_{\textit{intra}}^{*}=\max_{c\in\{1..C\}}\mathbb{E}[\cos(\mathbf{x}_{\textit{ID}},P_{*,c})]$, $\mu_{\textit{inter}}^{*}=\mathbb{E}[cos(\mathbf{x}_{\textit{ID}},I_{c_{j}})],\forall c_{j}\in\{1..C\}\setminus\{c_{\textit{ID}}\}$, and $C_{\textit{ID}}$ is the ground-truth class for the input $\mathbf{x}_{\textit{ID}}$.
With assumption 4, for OOD samples, we have
$\mathbb{E}\left[\max_{c\in\{1..C\}}\sigma_{c}^{\textit{image}}(\mathbf{x}_{\textit{OOD}})\right]=\mathbb{E}\left[\max_{c\in\{1..C\}}\sigma_{c}^{\textit{text}}(\mathbf{x}_{\textit{OOD}})\right]=\frac{1}{C}\,.$ (14)
Then we have
$\Delta_{\textit{image}}=\mathbb{E}\left[\max_{c\in\{1..C\}}\sigma_{c}^{\textit{image}}(\mathbf{x}_{i})\right]-\frac{1}{C}\,.$ (15)
Similarly, we have
$\Delta_{\textit{text}}=\mathbb{E}\left[\max_{c\in\{1..C\}}\sigma_{c}^{\textit{text}}(\mathbf{x}_{i})\right]-\frac{1}{C}\,.$ (16)
Since $\mu_{\textit{intra}}^{\textit{image}}>\mu_{\textit{intra}}^{\textit{text}}$ (Assumption 2), we have
$\Delta_{\textit{image}}$ $>\Delta_{\textit{text}}\,.$ (17)
By simple algebra, we have
$\mathbb{E}\left[\frac{\max_{i\in\{1..C\}}\sigma_{i}^{\textit{image}}(\mathbf{x}_{\textit{ID}})+\max_{i\in\{1..C\}}\sigma_{i}^{\textit{text}}(\mathbf{x}_{\textit{ID}})}{2}\right$. (18)
$\left.-\frac{\max_{i\in\{1..C\}}\sigma_{i}^{\textit{image}}(\mathbf{x}_{\textit{OOD}})+\max_{i\in\{1..C\}}\sigma_{i}^{\textit{text}}(\mathbf{x}_{\textit{OOD}})}{2}\right]$
$\geq$ $\mathbb{E}\left[\max_{i\in\{1..C\}}\sigma_{i}^{\textit{text}}(\mathbf{x}_{\textit{ID}})-\max_{i\in\{1..C\}}\sigma_{i}^{\textit{text}}(\mathbf{x}_{\textit{OOD}})\right]\,$,